OpenStudy (anonymous):

Help!! Definate Medal!!! Which lines are perpendicular to 5x + 4y = 8? Choose all answers that are correct. 4y=-5x+4 4y= 5x + 12 5y = 4x + 15 5y = –4x – 15 My answer is 4y=-5x+4.. what do you think

OpenStudy (anonymous):

@hero @Nnesha

OpenStudy (mrnood):

first write the equation as y = mx +b the value m is the slope of th eoriginal line all perpendicular lines will have slope = -1/m

OpenStudy (mrnood):

btw - I think your answer is incorrect...

OpenStudy (anonymous):

y = mx +b=4y=-5x+8

OpenStudy (mrnood):

you need to show some work - otherwise you can just guess all 4 until I say... what is the slope of the original line? then - what is the slope of any perpendicular line..?

OpenStudy (mrnood):

4y=-5x+8 so y= ??

OpenStudy (anonymous):

No I wasn't guessing all 4 -.-, you asked me to rewrite the equation in the formula y = mx +b so i did and i wanted you to take me from there.

OpenStudy (anonymous):

cause im kinda confused

OpenStudy (anonymous):

y=−5/4+2

OpenStudy (anonymous):

Slope is −5/4

OpenStudy (anonymous):

Y = 2 @MrNood

OpenStudy (anonymous):

but im confused from there..:(

OpenStudy (anonymous):

@Hero help please

OpenStudy (mrnood):

OK so slope of original line is -5/4 and so the slope of any perpendicular line is -1/m -1 /(-5/4) = 4/5 so which of the answer options has a slope of 4/5? (hint - the offset value i.e. the constant number does not affect the slope or answer....)